OpenStudy (anonymous):

Simplify: 3sqrt of -8 + 5 sqrt of -18

OpenStudy (calculusxy):

\[3\sqrt{-8} + 5\sqrt{-18}\] Is it this? Or if not, then can you please draw the problem?

OpenStudy (anonymous):

yes it is!

OpenStudy (calculusxy):

Well this seems to have no solution b/c you can't find the square root of a negative radicand.

OpenStudy (anonymous):

it would be an imaginary number right?

OpenStudy (solomonzelman):

Hello

OpenStudy (anonymous):

hello!

OpenStudy (calculusxy):

I am not too sure about imaginary numbers but i think that @SolomonZelman can help

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3\sqrt{-8}+5\sqrt{-18} }\) I would start from just getting the \(i\)'s out \(\large\color{black}{ \displaystyle 3i\sqrt{8}+5i\sqrt{18} }\)

OpenStudy (anonymous):

ok!

OpenStudy (solomonzelman):

then, I can re-write that as: \(\large\color{black}{ \displaystyle 3i\cdot\sqrt{4}\cdot \sqrt{2}+5i\cdot \sqrt{9}\cdot \sqrt{2} }\)

OpenStudy (solomonzelman):

can you finish from here?

OpenStudy (anonymous):

yes I can! @SolomonZelman , thank you!!!

OpenStudy (solomonzelman):

yw